OpenStudy (esraa):

A direct current power supply is connected to a circuit containing one resistor, and 1.5 A is drawn. If a second identical resistor is connected to the power supply in parallel with the first, how much current is drawn from the battery?

OpenStudy (anonymous):

V = IR resistors add in series, so it would be V = I (R1+R2) R1= R2, so V = 2RI you have two equations, and 2 unknowns, start eliminating :)

OpenStudy (esraa):

thnx =)

OpenStudy (anonymous):

no problem, good luck :)

OpenStudy (anonymous):

actually, you dont have to eliminate anything. V are the same for both cases. If you double R, then you obviously 1/2 I. My bad if i confused you, didn't read all, but yea. In most cases, you would have to eliminate. Good luck

OpenStudy (esraa):

Don'y worry this was already noticed, but thnx anyway :D

OpenStudy (anonymous):

The second resistor is in parallel. Did we correct that from the initial answer?

OpenStudy (anonymous):

my bad, parallel resistors add in 1/R + 1/R. Not R+R. Didn't read nicely again :[ but yea, same method, just that you add them 1/ R(equivalent) = 1/R+ 1/R

OpenStudy (esraa):

It won't make any difference in the final ans, isn't it? (I tried it and just want to confirm if this is right)

OpenStudy (anonymous):

it should actually.

OpenStudy (esraa):

Ok now I'm confused

OpenStudy (anonymous):

your current should double up I believe

OpenStudy (anonymous):

Since the resistors have the same value, the equivalent resistance is nice and easy \[R^* = {RR \over R+R} ={R \over 2}\] From Ohm's Law\[I = {V \over R}\] We observe that relationship between current and resistance is an inverse one. \[I \alpha {1 \over R}\] If we halve resistance, what must the current do such that voltage remains constant?

OpenStudy (anonymous):

V = I Req. = I R/2.

OpenStudy (esraa):

The current must e doubled so it doesn't make a difference if the resistors are in series or parallel, isn't it?

OpenStudy (anonymous):

No. The current will be halved if we hook the resistors up in series. \[R^*_{series} = R + R = 2R\] \[I = {V \over 2R}\] In parallel, \[I = {2 V \over R}\]

OpenStudy (anonymous):

yes, there is a huge difference! sorry for confusing you. Think of it like this. Current is like water, voltage is like pressure. If you put 2 claims in 1 hose, block the same pathway twice, you would flow extremely slow right? Now imagine the situation where you're checking in for a concert event. Would you wait longer if they check everyone twice by 1 gate, or when they have multiple gates to check multiple people once :)?

OpenStudy (esraa):

@eashmore thnx alot @vizkosity, it's ok now I got it =) Thnx again to both of u :D